OpenStudy (anonymous):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (0, 4) and (-2, 8) (3, 7) and (-1, -5) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Do you know how to find the slope?

OpenStudy (anonymous):

no :/

jimthompson5910 (jim_thompson5910):

Alright, I'll do one pair and hopefully you can use it to do the other pair (if not, let me know) Two generic points that define a straight line are (x1, y1) and (x2, y2) In this case, (x1, y1) = (0, 4), which means x1=0 and y1= 4 and (x2, y2) = (-2, 8), which means x2=-2 and y2= 8 So x1=0, y1= 4, x2=-2, and y2= 8. ------------------- Now use the slope formula m = (y2-y1)/(x2-x1) m = ( 8- 4)/(-2-0) m = 4/(-2) m = -2 So the slope of the line that passes through the two points (0, 4) and (-2, 8) is -2

jimthompson5910 (jim_thompson5910):

let me know if that helps or not, thanks

OpenStudy (anonymous):

The slope is 3

OpenStudy (anonymous):

For the other coordinates

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

so how do the slopes of -2 and 3 compare

OpenStudy (anonymous):

is it A?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

the answer is A

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

sure thing